# A unified convention for achievement positional games

### v1

@@ Abstract
We introduce achievement positional games, a convention for positional games which encompasses the Maker-Maker and Maker-Breaker conventions. We consider two hypergraphs, one red and one blue, on the same vertex set. Two players, Left and Right, take turns picking a previously unpicked vertex. Whoever first fills an edge of their color, blue for Left or red for Right, wins the game (draws are possible). We establish general properties of such games. In particular, we show that a lot of principles which hold for Maker-Maker games generalize to achievement positional games. We also study the algorithmic complexity of deciding whether Left has a winning strategy as first player when all blue edges have size at mot $p$ and all red edges have size at most $q$. This problem is in P for $p,q\leq 2$, but it is NP-hard for $p\geq 3$ and $q=2$, coNP-complete for $p=2$ and $q\geq 3$, and PSPACE-complete for $p,q\geq 3$. A consequence of this last result is that, in the Maker-Maker convention, deciding whether the first player has a winning strategy on a hypergraph of rank 4 after one round of (non-optimal) play is PSPACE-complete.
@@ 1 Introduction
Positional games. Positional games have been introduced by Hales and Jewett [HJ63] and later popularized by Erdős and Selfridge [ES73]. The game board is a hypergraph $H=(V,E)$, where $V$ is the vertex set and $E\subseteq 2^{V}$ is the edge set. Two players take turns picking a previously unpicked vertex of the hypergraph, and the result of the game is defined by one of several possible conventions. The two most popular conventions are called Maker-Maker and Maker-Breaker. As they revolve around trying to fill an edge i.e. pick all the vertices of some edge, they are often referred to as “achievement games”. In the Maker-Maker convention, whoever fills an edge first wins (draws are possible), whereas in the Maker-Breaker convention, Maker aims at filling an edge while Breaker aims at preventing him from doing so (no draw is possible). These games have counterparts in the form of “avoidance games”. In the Avoider-Avoider convention, whoever fills an edge first loses, whereas in the Avoider-Enforcer convention, Avoider aims at avoiding filling an edge while Enforcer aims at forcing her to do so. For all these conventions, the central question is to find out what the result of the game is with optimal play (who wins, or is it a draw). The study of positional games mainly consists, for various conventions and classes of hypergraphs, in finding necessary and/or sufficient conditions for such player to have a winning strategy (on the number of edges, the size of the edges, the structure of the hypergraph…), and determining the complexity of the associated algorithmic problems.
@@
The Maker-Maker convention was the first one to be introduced, in 1963 by Hales and Jewett [HJ63]. The game of Tic-Tac-Toe is a famous example. As a simple strategy-stealing argument [HJ63] shows that the second player cannot have a winning strategy, the question is whether the given hypergraph $H$ is a first player win or a draw with optimal play. This decision problem is trivially tractable for hypergraphs of rank 2 i.e. whose edges have size at most 2, but it is PSPACE-complete for hypergraphs that are 6-uniform i.e. whose edges have size exactly 6 (by a combination of [Koe25] and [Bys04]). Maker-Maker games are notoriously difficult to handle. They are strong games, meaning that both players must manage offense and defense at the same time, by trying to fill an edge while also preventing the opponent from doing so first. As such, it is unclear which player should benefit even from some basic hypergraph operations such as adding an edge (see the extra edge paradox [Bec08]).
@@
The Maker-Breaker convention was introduced for that reason, in 1978 by Chvátal and Erdős [CE78]. It is by far the most studied, as it presents some convenient additional properties compared to the Maker-Maker convention thanks to the players having complementary goals. The most crucial property is subhypergraph monotonicity: if Maker has a winning strategy on a subhypergraph of $H$, then he may disregard the rest of the hypergraph and apply that strategy to win on $H$, as he has no defense to take care of. The problem of deciding which player has a winning strategy when, say, Maker starts, is tractable for hypergraphs of rank 3 [GGS22] but PSPACE-complete for 5-uniform hypergraphs [Koe25] (a very recent improvement on the previously known result for 6-uniform hypergraphs [RW21]). Consider the board game Hex for instance, where two players take turns placing tokens of their color to try and connect opposite sides of the board. It is well-known [Gar59] that a draw is technically impossible, and that successfully connecting one’s sides of the board is equivalent to blocking the opponent from connecting theirs. As such, Hex is often presented as a Maker-Breaker game, but it only is so by theorem. By nature, it actually is an achievement game which does not fall under either Maker-Maker or Maker-Breaker conventions, as the first player to fill an edge wins but the players have different edges to fill (corresponding here to “horizontal paths” and “vertical path” respectively).
@@
Unified achievement games. We introduce achievement positional games. Such a game is a triple $\mathcal{G}=(V,E_{L},E_{R})$, where $(V,E_{L})$ and $(V,E_{R})$ are hypergraphs which we see as having blue edges and red edges respectively. There are two players, taking turns picking a previously unpicked vertex: Left aims at filling a blue edge, while Right aims at filling a red edge. Whoever reaches their goal first wins the game, or we get a draw if this never happens. Achievement positional games include all Maker-Maker and Maker-Breaker games. Indeed, Maker-Maker games correspond to the case $E_{L}=E_{R}$, while Maker-Breaker games correspond to the case $E_{R}=\varnothing$ when identifying Maker with Left and Breaker with Right and renaming Breaker wins as draws (although there is another way to embed Maker-Breaker games into achievement positional games, as we explain later in this paragraph). We note that this idea of “edge-partizan” positional games appears briefly in [BDD+24], albeit in the context of scoring games, when showing that the scoring versions of the Maker-Maker and Maker-Breaker conventions belong to Milnor’s universe.
@@
Objectives and results. The aim of this paper is twofold. We first establish elementary properties of achievement positional games in general. In particular, we look at some general principles which hold in the Maker-Maker convention, such as strategy stealing for instance, to see if they generalize to achievement positional games. For all those that we consider, we show that this is indeed the case, emphasizing the fact that most properties of Maker-Maker games come from their “achievement” nature rather than symmetry. We also define the outcome of an achievement positional game as the result with optimal play when Left starts and when Right starts. We list all existing outcomes, as well as all possibilities for the outcome of a disjoint union of two games depending on their individual outcomes. Our second objective is the study of the algorithmic complexity of the game, depending on the size of both players’ edges. The problem we consider is the existence of a winning strategy for Left as first player. We get results for almost all edge sizes, which are summed up in Table 1. As a corollary, we also show that deciding whether the next player has a winning strategy for the Maker-Maker game on a hypergraph of rank 4 after one round of (non-optimal) play is PSPACE-complete.
$q$ $p$ 0 , 1 2 3 4 5+
0 , 1 LSPACE [trivial] LSPACE [RW20] P [GGS22] open PSPACE-c [Koe25]
2 LSPACE [trivial] P [Th. 4.1] NP-hard [Th. 4.5] NP-hard [Th. 4.5] PSPACE-c [Koe25]
3+ LSPACE [trivial] coNP-c [Th. 4.3] PSPACE-c [Th. 4.6] PSPACE-c [Th. 4.6] PSPACE-c [Koe25]
Table 1: Algorithmic complexity of deciding whether Left has a winning strategy as first player, for blue edges of size at most $p$ and red edges of size at most $q$.
@@ 2.1 Definitions
In this paper, a hypergraph is a pair $(V,E)$ where $V$ is a finite vertex set and $E\subseteq 2^{V}\setminus\{\varnothing\}$ is the edge set. An achievement positional game is a triple $\mathcal{G}=(V,E_{L},E_{R})$ where $(V,E_{L})$ and $(V,E_{R})$ are hypergraphs. The elements of $E_{L}$ are called blue edges, whereas the elements of $E_{R}$ are called red edges. Two players, Left and Right, take turns picking a vertex in $V$ that has not been picked before. We say a player fills an edge if that player has picked all the vertices of that edge. The blue and red edges can be seen as the winning sets of Left and Right respectively, so that the result of the game is determined as follows:
@@
The player who starts the game may be either Left or Right. Therefore, when talking about winning strategies (i.e. strategies that guarantee a win) or non-losing strategies (i.e. strategies that guarantee a draw or a win), we will always specify which player is assumed to start the game. For instance, we may say that “Left has a winning strategy on $\mathcal{G}$ as first player”.
@@ 2.2 Updating the edges
If Right had picked $\gamma_{1}$ or $\gamma_{2}$ instead of $\beta_{1}$, then Left would have won the same way. If Right had picked $\beta_{2}$, $\gamma_{3}$ or $\gamma_{4}$ instead of $\beta_{1}$, then Left would have picked $\beta_{1}$ and won in similar fashion. All in all, Left has a winning strategy (in three moves) as first player on the blue butterfly. Obviously, if Right starts then optimal play leads to a draw with Right picking $\alpha$. The butterfly, in its red or blue version, will be useful for constructing games such as hardness gadgets.
@@ 2.3 Decision problems
It actually suffices to address one of these combinations. First of all, we need only consider strategies for Left, as the question for Right is the same up to swapping $E_{L}$ and $p$ with $E_{R}$ and $q$ respectively. Moreover, the problem of non-losing strategies is the complement of the problem of winning strategies. Finally, we can assume that Left starts, as this problem admits a linear-time reduction to the case where Right starts. Indeed, Left has a winning strategy on $\mathcal{G}=(V,E_{L},E_{R})$ as second player if and only if, for all $u\in V$, we have $\{u\}\not\in E_{R}$ and Left has a winning strategy on $\mathcal{G}^{u}$ as first player (note that $\mathcal{G}^{u}$ still has the property that all blue edges have size at most $p$ and all red edges have size at most $q$). All in all, we introduce the following decision problem.
@@
Output : T if Left has a winning strategy on $\mathcal{G}$ as first player, F otherwise.
We can start by noticing that, for all $p,q$, AchievementPos($p$,$q$) is in PSPACE. This is due to the fact that the game cannot last more than $|V|$ moves, so even a brute-force algorithm only needs a space $O(|V|)$ to determine the winner (see [Sch78, Lemma 2.2] for more details).
This problem encompasses the decision problems associated to the Maker-Maker and Maker-Breaker conventions.
@@
Output : T if Maker has a winning strategy as first player for the Maker-Breaker game on $H$, F otherwise.
Note that MakerMaker($k$) is the subproblem of AchievementPos($k$,$k$) where $E_{L}=E_{R}$. Moreover, MakerBreaker($k$) is exactly equivalent to AchievementPos($k$,00). Indeed, in AchievementPos($k$,00), we have $E_{R}=\varnothing$ and Right is therefore reduced to a “Breaker” role, the only difference being that a Breaker win is considered a draw. Similarly, AchievementPos($k$,$1$) is also equivalent to MakerBreaker($k$). Indeed, the case where $E_{R}$ has at least two elements of size 1 is trivial, and the case where $E_{R}$ has exactly one element of size 1 reduces to AchievementPos($k$,00) after one round of optimal play.
Let us recall known complexity results for the Maker-Maker and Maker-Breaker conventions. It is easy to see that MakerMaker($2$) and MakerBreaker($2$) are in LSPACE [RW20]. Moreover, MakerBreaker($3$) is in P [GGS22]. As for hardness results, it has recently been shown that MakerBreaker($5$) is PSPACE-complete [Koe25], improving on the previously known result for MakerBreaker($6$) [RW21]. Using a simple argument by Byskov [Bys04], this implies that MakerMaker($6$) is PSPACE-complete. The algorithmic complexity of MakerMaker($3$), MakerMaker($4$), MakerMaker($5$) and MakerBreaker($4$) is currently open.
@@
AchievementPos($k$,00) and AchievementPos($k$,$1$) are in LSPACE for $k\geq 2$, in P for $k=3$, but are PSPACE-complete for $k\geq 5$. Moreover, AchievementPos(00,$k$) and AchievementPos($1$,$k$) are in LSPACE for all $k$.
@@
It is clear that AchievementPos(00,$k$) and AchievementPos($1$,$k$) are in LSPACE for all $k$: indeed, if all the red edges have size 1, then Left has a winning strategy as first player if and only if $E_{R}\neq\varnothing$. The results on AchievementPos($k$,00) and AchievementPos($k$,$1$) are a consequence of the known results (listed above) about MakerBreaker($k$) and the fact that this problem is equivalent to AchievementPos($k$,00) and AchievementPos($k$,$1$). ∎
@@ 2.4 Outcome
  • If Left has a winning (resp. non-losing) strategy as first player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as first player on $\mathcal{G}_{u}$.
  • If Left has a winning (resp. non-losing) strategy as second player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as second player on $\mathcal{G}_{u}$.
@@
Let $\mathcal{G}$ be an achievement positional game. If Left has a winning (resp. non-losing) strategy as second player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as first player on $\mathcal{G}$.
@@
Suppose that Left has a winning (resp. non-losing) strategy as second player on $\mathcal{G}$. Playing first on $\mathcal{G}$, Left can pick an arbitrary vertex $u$. We may assume that Left does not instantly win the game doing so. Now, on the updated game $\mathcal{G}_{u}$, Left has a winning (resp. non-losing) strategy as second player by Lemma 2.2. ∎

### v2

@@ Abstract
We introduce achievement positional games, a convention for positional games which encompasses the Maker-Maker and Maker-Breaker conventions. We consider two hypergraphs, one red and one blue, on the same vertex set. Two players, Left and Right, take turns picking a previously unpicked vertex. Whoever first fills an edge of their color, blue for Left or red for Right, wins the game (draws are possible). We establish general properties of such games. In particular, we show that a lot of principles which hold for Maker-Maker games generalize to achievement positional games. We also study the algorithmic complexity of deciding whether Left has a winning strategy as the first player when blue edges and red edges have respective sizes at most $p$ and $q$. This problem is in P for $p,q\leq 2$, but it is NP-hard for $p\geq 3$ and $q=2$, coNP-complete for $p=2$ and $q\geq 3$, and PSPACE-complete for $p,q\geq 3$ even when the 3-edges are the same for both colors. That last result has an interesting consequence on the Maker-Maker convention: for 3-uniform hypergraphs, which is the only case whose complexity is currently open (for starting positions of the game), we show PSPACE-completeness for positions obtained after one round of play.
@@ 1 Introduction
Positional games. Positional games have been introduced by Hales and Jewett [13] and later popularized by Erdős and Selfridge [6]. The game board is a hypergraph $H=(V,E)$, where $V$ is the vertex set and $E\subseteq 2^{V}$ is the edge set. Two players take turns picking a previously unpicked vertex of the hypergraph, and the result of the game is defined by one of several possible conventions. The two most popular conventions are called Maker-Maker and Maker-Breaker. As they revolve around trying to fill an edge i.e. pick all the vertices of some edge, they are often referred to as “achievement games”. In the Maker-Maker convention, whoever fills an edge first wins (draws are possible), whereas in the Maker-Breaker convention, Maker aims at filling an edge while Breaker aims at preventing him from doing so (no draw is possible). These games have counterparts in the form of “avoidance games”. In the Avoider-Avoider convention, whoever fills an edge first loses, whereas in the Avoider-Enforcer convention, Avoider aims at avoiding filling an edge while Enforcer aims at forcing her to do so. For all these conventions, the central question is to find out what the result of the game is with optimal play (who wins, or is it a draw). The study of positional games mainly consists, for various conventions and classes of hypergraphs, in finding necessary and/or sufficient conditions for such player to have a winning strategy (on the number of edges, the size of the edges, the structure of the hypergraph…), and determining the complexity of the associated algorithmic problems.
@@
The Maker-Maker convention was the first one to be introduced, in 1963 by Hales and Jewett [13]. The game of Tic-Tac-Toe is a famous example. As a simple strategy-stealing argument [13] shows that the second player cannot have a winning strategy, the question is whether the given hypergraph $H$ is a first player win or a draw with optimal play. This decision problem is trivially tractable for hypergraphs of rank 2 i.e. whose edges have size at most 2, but it is PSPACE-complete for hypergraphs that are 4-uniform i.e. whose edges have size exactly 4 [10, 11]. Maker-Maker games are notoriously difficult to handle. They are strong games, meaning that both players must manage offense and defense at the same time, by trying to fill an edge while also preventing the opponent from doing so first. As such, it is unclear which player should benefit even from some basic hypergraph operations such as adding an edge (see the extra edge paradox [2]).
@@
The Maker-Breaker convention was introduced for that reason, in 1978 by Chvátal and Erdős [5]. It is by far the most studied, as it presents some convenient additional properties compared to the Maker-Maker convention thanks to the players having complementary goals. The most crucial property is subhypergraph monotonicity: if Maker has a winning strategy on a subhypergraph of $H$, then he may disregard the rest of the hypergraph and apply that strategy to win on $H$, as he has no defense to take care of. The problem of deciding which player has a winning strategy when, say, Maker starts, is tractable for hypergraphs of rank 3 [8] but PSPACE-complete for 4-uniform hypergraphs [11] (a very recent improvement on the previously known results for 6-uniform hypergraphs [19] and 5-uniform hypergraphs [15]). Consider the board game Hex for instance, where two players take turns placing tokens of their color to try and connect opposite sides of the board. It is well-known [12] that a draw is technically impossible, and that successfully connecting one’s sides of the board is equivalent to blocking the opponent from connecting theirs. As such, Hex is often presented as a Maker-Breaker game, but it only is so by theorem. By nature, it actually is an achievement game which does not fall under either Maker-Maker or Maker-Breaker conventions, as the first player to fill an edge wins but the players have different edges to fill (corresponding here to “horizontal paths” and “vertical paths” respectively).
@@
Unified achievement games. We introduce achievement positional games. Such a game is a triple $\mathcal{G}=(V,E_{L},E_{R})$, where $(V,E_{L})$ and $(V,E_{R})$ are hypergraphs which we see as having blue edges and red edges respectively. There are two players, Left and Right, taking turns picking a previously unpicked vertex: Left aims at filling a blue edge, while Right aims at filling a red edge. Whoever reaches their goal first wins the game, or we get a draw if this never happens. Achievement positional games include all Maker-Maker and Maker-Breaker games. Indeed, Maker-Maker games correspond to the case $E_{L}=E_{R}$, while Maker-Breaker games correspond to the case $E_{R}=\varnothing$ when identifying Maker with Left and Breaker with Right and renaming Breaker wins as draws (although there is another way to embed Maker-Breaker games into achievement positional games, as we explain later in this paragraph). We note that this idea of “edge-partizan” positional games appears briefly in [1], albeit in the context of scoring games, when showing that the scoring versions of the Maker-Maker and Maker-Breaker conventions belong to Milnor’s universe.
@@
Objectives and results. The aim of this paper is twofold. We first establish elementary properties of achievement positional games in general. In particular, we look at some general principles which hold in the Maker-Maker convention, such as strategy stealing for instance, to see if they generalize to achievement positional games. For all those that we consider, we show that this is indeed the case, emphasizing the fact that most properties of Maker-Maker games come from their “achievement” nature rather than symmetry. We also define the outcome of an achievement positional game as the result with optimal play when Left starts and when Right starts. We list all existing outcomes, as well as all possibilities for the outcome of a disjoint union of two games depending on their individual outcomes. Our second objective is the study of the algorithmic complexity of the game, depending on the size of both players’ edges. The problem we consider is the existence of a winning strategy for Left as the first player. For some edge sizes, the algorithmic complexity of this problem ensues from previous results on positional games. We obtain results for all the remaining edge sizes, which can be seen in Table 1. A corollary of our work is that deciding whether the first player has a winning strategy for the Maker-Maker game on a 3-uniform hypergraph after one round of (non-optimal) play is PSPACE-complete, which is the first general complexity result on the Maker-Maker convention for edges of size 3.
$q$ pp 0 , 1 2 3 4 5+
0 , 1 LSPACE [trivial] LSPACE [18] P [8] PSPACE-c [11] PSPACE-c [15]
2 LSPACE [trivial] P [Th. 4.1] NP-hard [Th. 4.6] PSPACE-c [11] PSPACE-c [15]
3+ LSPACE [trivial] coNP-c [Th. 4.4] PSPACE-c [Th. 4.7] PSPACE-c [11] PSPACE-c [15]
Table 1: Algorithmic complexity of deciding whether Left has a winning strategy as the first player, for blue edges of size at most $p$ and red edges of size at most $q$. Yellow cells correspond to this paper’s main results.
We note that an extended abstract of this work was published at EuroComb’25 [9], however the PSPACE-hardness proof for the case $(p,q)=(3,3)$ was different and did not imply PSPACE-hardness for intermediate positions of the Maker-Maker game on 3-uniform hypergraphs, whereas the proof presented in this paper does. On the other hand, a recent paper has built on this extended abstract to improve on one of our results: namely, the case $(p,q)=(3,2)$, which Theorem 4.6 proves to be NP-hard, has been shown to be PSPACE-complete [10].
@@ 2.1 Definitions
In this paper, a hypergraph is a pair $(V,E)$ where $V$ is a finite vertex set and $E\subseteq 2^{V}\setminus\{\varnothing\}$ is the edge set. An edge of size $k$ may be called a $k$-edge. An achievement positional game is a triple $\mathcal{G}=(V,E_{L},E_{R})$ where $(V,E_{L})$ and $(V,E_{R})$ are hypergraphs. The elements of $E_{L}$ are called blue edges, whereas the elements of $E_{R}$ are called red edges. Two players, Left and Right, take turns picking a vertex in $V$ that has not been picked before. We say a player fills an edge if that player has picked all the vertices of that edge. The blue and red edges can be seen as the winning sets of Left and Right respectively, so that the result of the game is determined as follows:
@@
The player who starts the game may be either Left or Right. Therefore, when talking about winning strategies (i.e. strategies that guarantee a win) or non-losing strategies (i.e. strategies that guarantee a draw or a win), we will always specify which player is assumed to start the game. For instance, we may say that “Left has a winning strategy on $\mathcal{G}$ as the first player”.
@@ 2.2 Updating the edges
If Right had picked $\gamma_{1}$ or $\gamma_{2}$ instead of $\beta_{1}$, then Left would have won the same way. If Right had picked $\beta_{2}$, $\gamma_{3}$ or $\gamma_{4}$ instead of $\beta_{1}$, then Left would have picked $\beta_{1}$ and won in similar fashion. All in all, Left has a winning strategy (in three moves) as the first player on the blue butterfly. Obviously, if Right starts, then optimal play leads to a draw with Right picking $\alpha$. The butterfly, in its red or blue version, will be useful for constructing games such as hardness gadgets.
@@ 2.3 Decision problems
It actually suffices to address one of these combinations. First of all, we need only consider strategies for Left, as the question for Right is the same up to swapping $E_{L}$ and $p$ with $E_{R}$ and $q$ respectively. Moreover, the problem of non-losing strategies is the complement of the problem of winning strategies. Finally, we can assume that Left starts, as this problem admits a linear-time reduction to the case where Right starts. Indeed, Left has a winning strategy on $\mathcal{G}=(V,E_{L},E_{R})$ as the second player if and only if, for all $u\in V$, we have $\{u\}\not\in E_{R}$ and Left has a winning strategy on $\mathcal{G}^{u}$ as the first player (note that the maximum size of a blue (resp. red) edge in of $\mathcal{G}^{u}$ is at most the maximum size of a blue (resp. red) edge in $\mathcal{G}$). All in all, we introduce the following decision problem.
@@
Output : T if Left has a winning strategy on $\mathcal{G}$ as the first player, F otherwise.
Proposition 2.1.
AchievementPos($p$,$q$) is in PSPACE for all $p,q$.
Proof.
Since the length of the game is bounded by the number of vertices, AchievementPos($p$,$q$) belongs to the class AP of problems decidable in polynomial time by an alternating Turing machine. The result ensues from the fact that AP=PSPACE [4]. ∎
This problem encompasses the decision problems associated with the Maker-Maker and Maker-Breaker conventions.
@@
Output : T if Maker has a winning strategy as the first player for the Maker-Breaker game on $H$, F otherwise.
Note that MakerMaker($k$) is the subproblem of AchievementPos($k$,$k$) where $E_{L}=E_{R}$. Moreover, MakerBreaker($k$) is equivalent to AchievementPos(kk,0). Indeed, in AchievementPos(kk,0), we have $E_{R}=\varnothing$ and Right is therefore reduced to a “Breaker” role, the only difference being that a Breaker win is considered a draw. Similarly, AchievementPos(kk,11) is also equivalent to MakerBreaker($k$). Indeed, the case where there are at least two 1-edges in $E_{R}$ is trivial, and the case where there is exactly one 1-edge in $E_{R}$ reduces to AchievementPos(kk,0) after one round of optimal play.
Let us recall known complexity results for the Maker-Maker and Maker-Breaker conventions. It is easy to see that MakerMaker($2$) and MakerBreaker($2$) are in LSPACE [18]. Moreover, MakerBreaker($3$) is in P [8]. As for hardness results, it has recently been shown that MakerBreaker($4$) is PSPACE-complete [11], improving on the previously known result for MakerBreaker($5$) [15], and that MakerMaker($4$) is PSPACE-complete [10]. The algorithmic complexity of MakerMaker($3$) is currently open.
@@
AchievementPos(pp,0) and AchievementPos(pp,11) are in LSPACE for all $p\geq 2$, and in P for $p=3$, but are PSPACE-complete for all $p\geq 4$. Moreover, AchievementPos(0,$q$) and AchievementPos(11,$q$) are in LSPACE for all $q$.
@@
It is clear that AchievementPos(0,$q$) and AchievementPos(11,$q$) are in LSPACE for all $q$: indeed, if all the red edges have size $1$, then Left has a winning strategy as the first player if and only if $E_{R}\neq\varnothing$. The results on AchievementPos(pp,0) and AchievementPos(pp,11) are a consequence of the known results (listed above) about MakerBreaker($p$), which is equivalent to AchievementPos(pp,0) and AchievementPos(pp,11). ∎
@@ 2.4 Outcome
  • If Left has a winning (resp. non-losing) strategy as the first player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as the first player on $\mathcal{G}_{u}$.
  • If Left has a winning (resp. non-losing) strategy as the second player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as the second player on $\mathcal{G}_{u}$.
@@
Let $\mathcal{G}$ be an achievement positional game. If Left has a winning (resp. non-losing) strategy as the second player on $\mathcal{G}$, then Left has a winning (resp. non-losing) strategy as the first player on $\mathcal{G}$.
@@
Suppose that Left has a winning (resp. non-losing) strategy as the second player on $\mathcal{G}$. Playing first on $\mathcal{G}$, Left can pick an arbitrary vertex $u$. We may assume that Left does not instantly win the game doing so. Now, on the updated game $\mathcal{G}_{u}$, Left has a winning (resp. non-losing) strategy as the second player by Lemma 2.3. ∎